OpenStudy (anonymous):

Which choice shoes the sum of the two vectors <-2,0> and <4,-5> written as and ordered pair A. <4,-5> B. <2,-5> C. <-2,0> D. <-6.5>

OpenStudy (johnweldon1993):

So you would add the 'x' components and then add the y'components seperately <-2,0> + <4,-5> = <-2 + 4 , 0 - 5>

OpenStudy (anonymous):

Thank you!! do you think you could help me with another/

OpenStudy (anonymous):

I can :)

OpenStudy (anonymous):

The ordered pairs <-8,2> and <5,-6> represent two vectors. What is the magnitude of the resultant?

OpenStudy (anonymous):

Resultant vector: <–8, 2> + <5, –6> = <-8+5, 2+(-6)> = <-3, -4> Magnitude of resultant vector = sqrt((-3)^2+(-4)^2) = sqrt(9+16)=sqrt(25) = 5

OpenStudy (anonymous):

the simple answer is 5

OpenStudy (anonymous):

that was a choice, i just have a few more, im really bad at geometry haha

OpenStudy (anonymous):

Okay! haha

OpenStudy (anonymous):

The ordered pairs <5,2> and <3,-2> represent two vectors. What is the magnitude of the resultant?

OpenStudy (anonymous):

that one is 6

OpenStudy (johnweldon1993):

Are you sure @gurrilabill ?

OpenStudy (anonymous):

do you disagree?

OpenStudy (johnweldon1993):

<5 + 3 , 2 - 2> = resultant = <8 , 0> \[\sqrt {(8)^2 + (0)^2} = \sqrt{64} = 8\]

OpenStudy (anonymous):

Ahhh, Sorry about that abbi, Thanks john!

OpenStudy (anonymous):

aw its okay, thank you for helping me though!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

I have a few more D;

OpenStudy (anonymous):

Okay!!

OpenStudy (johnweldon1993):

Right...to find the 'x' value...you do x = 5 * cos(13) and to find the 'y' it would be y = 5 * sin(13) Which would arrive to answer choice C

OpenStudy (anonymous):

sorry guys, i know im asking a lot, i just have no clue what im doing

OpenStudy (anonymous):

Use the appropriate trig function, sin and cos, to get the x and y components. One for the x-compnonet and the other one for the y-component. For that triangle, the side measuring 5 is obviously the hypotenuse. Relative to the 13-deg angle, The x-compnent is the adjacent leg. Since you have the hypotenuse, and you are looking for the adjacent leg, you need the cosine. The y-compnent is the oppoite leg. The function that relates the opposite leg and the hypotenuse is the sine. cos 13 = x/5 sin 13 = y/5 x = 5*cos13 y = 5*sin13 The x and y numbers will come out positive as if they were in the first quadrant. Since your vector terminates in the third quadrant, just make the values of x and y negative. x = 4.87185... y = 1.124755...

OpenStudy (johnweldon1993):

Well you know it is definitely going south...so that gets rid of answers 2 and 3... Are you going west? No, it is pointing towards the east... so south ...east....looks like answer choice 4

OpenStudy (anonymous):

i dont know this one :/ @johnweldon1993 do you know?

OpenStudy (johnweldon1993):

Indeed I do....so it looks to be the person looking UP at the radar tower....elevated angle....so it would be A" The angle of elevation from the person to the radar tower

OpenStudy (anonymous):

im really sorry, just a few more?

OpenStudy (anonymous):

@johnweldon1993 you can take it from here, sorry abbi i have to get off for a few